# Eyes Wide Shut? Exploring the Visual Shortcomings of Multimodal LLMs

### v1

@@ 3.1 Visual Patterns in CLIP-blind Pairs
  OpenAI ViT-L-14 [43] ${}^{2}$ 427.6 75.5 13.3 13.3 20.0 20.0 13.3 53.3 20.0 6.7 13.3 19.3
[…]
SigLIP ViT-SO-14 [66] ${}^{2}$ 877.4 82.0 26.7 20.0 53.3 40.0 20.0 66.7 40.0 20.0 53.3 37.8
[…]
DFN ViT-H-14 [10] ${}^{2}$ 986.1 83.4 20.0 26.7 73.3 26.7 26.7 66.7 46.7 13.3 53.3 39.3
[…]
MetaCLIP ViT-L-14 [62] ${}^{2}$ 427.6 79.2 13.3 6.7 66.7 6.7 33.3 46.7 20.0 6.7 13.3 23.7
MetaCLIP ViT-H-14 [62] ${}^{2}$ 986.1 80.6 6.7 13.3 60.0 13.3 6.7 53.3 26.7 13.3 33.3 25.2
EVA01 ViT-g-14 [54] ${}^{2}$ 1136.4 78.5 6.7 26.7 40.0 6.7 13.3 66.7 13.3 13.3 20.0 23.0
EVA02 ViT-bigE-14+ [54] ${}^{2}$ 5044.9 82.0 13.3 20.0 66.7 26.7 26.7 66.7 26.7 20.0 33.3 33.3
@@ 4.3 Interleaved MoF
LLaVA ${}^{2}$ 256 5.5 81.8 50.0
[…]
LLaVA + I-MoF ${}^{2}$ 512 16.7(+10.7) 82.8 51.0
[…]
@@
We summarize the results in Table 3. Under the LLaVA setting, interleave MoF significantly enhances visual grounding, with a 10.7% increase observed in MMVP, without compromising the model’s ability to follow instructions. This experiment is replicated with the LLaVA-1.5 setting and under various image resolution settings, yielding similar enhancements in performance. We also evaluate on POPE [27] which is designed to test hallucination in visual grounding. Interleaved-MoF also shows consistent improvement against the original LLaVA models. Merely increasing the image resolution, and consequently, the number of tokens does not boost visual grounding capabilities. Instead, it is the interleaving of MoF that leads to improved performance in visual grounding tasks. We also evaluated Interleaved MoF on additional benchmarks such as MM-Bench [32] and GQA [21], finding that Interleaved MoF achieves similar performance on these benchmarks. Please refer to Appendix E for more results on these benchmarks.
@@ 6 Discussion
Acknowledgements. We thank Penghao Wu, Muzi Tao, Erik Jones, Michael Psenka, Daniel Yeh, Druv Pai for helpful discussions and feedback. We also thank Google Cloud and the TRC program for their support.
@@ B.2 Questions in MMVP Benchmark
We present more examples in MMVP at the end in Figures 101112. We also share the entire benchmark in the supplementary material.

### v2

@@ 3.1 Visual Patterns in CLIP-blind Pairs
  OpenAI ViT-L-14 [43] 2242 427.6 75.5 13.3 13.3 20.0 20.0 13.3 53.3 20.0 6.7 13.3 19.3
OpenAI ViT-L-14 [43] 3362 427.9 76.6 0.0 20.0 40.0 20.0 6.7 20.0 33.3 6.7 33.3 20.0
SigLIP ViT-SO-14 [66] 2242 877.4 82.0 26.7 20.0 53.3 40.0 20.0 66.7 40.0 20.0 53.3 37.8
SigLIP ViT-SO-14 [66] 3842 878.0 83.1 20.0 26.7 60.0 33.3 13.3 66.7 33.3 26.7 53.3 37.0
DFN ViT-H-14 [10] 2242 986.1 83.4 20.0 26.7 73.3 26.7 26.7 66.7 46.7 13.3 53.3 39.3
DFN ViT-H-14 [10] 3782 986.7 84.4 13.3 20.0 53.3 33.3 26.7 66.7 40.0 20.0 40.0 34.8
MetaCLIP ViT-L-14 [62] 2242 427.6 79.2 13.3 6.7 66.7 6.7 33.3 46.7 20.0 6.7 13.3 23.7
MetaCLIP ViT-H-14 [62] 2242 986.1 80.6 6.7 13.3 60.0 13.3 6.7 53.3 26.7 13.3 33.3 25.2
EVA01 ViT-g-14 [54] 2242 1136.4 78.5 6.7 26.7 40.0 6.7 13.3 66.7 13.3 13.3 20.0 23.0
EVA02 ViT-bigE-14+ [54] 2242 5044.9 82.0 13.3 20.0 66.7 26.7 26.7 66.7 26.7 20.0 33.3 33.3
@@ 4.3 Interleaved MoF
LLaVA 2242 256 5.5 81.8 50.0
LLaVA 3362 576 6.0 81.4 50.1
LLaVA + I-MoF 2242 512 16.7(+10.7) 82.8 51.0
LLaVA1.5 3362 576 24.7 84.7 85.9
LLaVA1.5 + I-MoF 2242 512 28.0(+3.3) 82.7 86.3
@@
We summarize the results in Table 3. Under the LLaVA setting, interleave MoF significantly enhances visual grounding, with a 10.7% increase observed in MMVP, without compromising the model’s ability to follow instructions. This experiment is replicated with the LLaVA-1.5 setting and under various image resolution settings, yielding similar enhancements in performance. We also evaluate on POPE [27] which is designed to test hallucination in visual grounding. Interleaved-MoF also shows consistent improvement against the original LLaVA models. Merely increasing the image resolution, and consequently, the number of tokens does not boost visual grounding capabilities. Instead, it is the interleaving of MoF between vision-only SSL models and VLM models that leads to improved performance in visual grounding tasks. We conduct more experiments using MAE or MoCoV3 as vision-only SSL models in I-MoF and show similar improvements in visual grounding tasks in Appenfix E.1. We also evaluated Interleaved MoF on additional benchmarks such as MM-Bench [32] and GQA [21], finding that Interleaved MoF achieves similar performance on these benchmarks. Please refer to Appendix E.2 for more results on these benchmarks.
@@ 6 Discussion
Acknowledgements. We thank Penghao Wu, Muzi Tao, Erik Jones, Michael Psenka, Daniel Yeh, Druv Pai, Chen Sun for helpful discussions and feedback. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise. This research is also supported by Intel, Google TRC program, the Google Cloud Research Credits program with the award GCP19980904, and an Amazon Research Award Fall 2023. The authors thank hyperbolic labs for supporting part of the experiments. All experiments and data processing were performed at NYU.
@@ B.2 Questions in MMVP Benchmark
We present more examples in MMVP at the end in Figures 101112.
@@ Appendix E More Benchmark Results
E.1 Different vision-only backbones
Here, we conduct extra experiments to study MoF involving MAE [18] or MoCoV3 [17] instead of DINOv2; See Table 6. In Table 6, we observe that with MAE/MoCov3, there is a consistent improvement in visual grounding ability, as shown in the MMVP and POPE benchmarks.
method SSL Model res #tokens MMVP POPE
LLaVA1.5 None 3362 576 24.7 85.9
LLaVA1.5 + I-MoF MoCov3 2242 512 26.7(+2.0) 86.1
LLaVA1.5 + I-MoF MAE 2242 512 27.3(+2.6) 86.1
LLaVA1.5 + I-MoF DINOv2 2242 512 28.0(+3.3) 86.3
Table 6: Results of Interleaved MoF with different vision-only SSL model
E.2 Scaling up to larger resolution
[…]
LLaVA1.5 3362 576 24.7 84.7 70.7 67.7 61.3 85.9 80.0 35.4
LLaVA1.5 + I-MoF 2242 512 28.0 82.7 73.3 61.6 55.3 86.3 77.3 33.5
LLaVA1.5 + I-MoF 3362 1152 31.3 81.8 73.3 65.4 58.7 86.7 79.3 34.6
[…]